A solution of vanadium(IV) ions is oxidized into vanadium(V) ions by reacting it with a standard solution of potassium hypochlorite. Use the following evidence to determine the concentration of the potassium hypochlorite solution:

Titration of 20.00 mL of a 0.0218 mol/L vanadium(IV) solution with a potassium hypochlorite solution.

Trial 1 2 3 4 5 6

Initial Buret Reading (mL) 0.00 6.50 12.60 19.33 25.51 31.57

Final Buret Reading (mL) 6.50 12.60 19.33 25.51 31.57 37.74


1
Expert's answer
2021-03-24T01:06:03-0400

final - initial titre values gave the following for trial 2-6

6.1 6.73 6.18 6.06 6.17

Average volume of hypochlorite = (6.1+6.73 +6.18+ 6.06 +6.17)/5 = 6.25


the mole ratio of hypochlorite to vanadium oxide is 1:2

Conc of hypochlorite = "\dfrac{1}{2}(\dfrac{0.0218 * 20}{6.25}) =0.035M"